OpenStudy (anonymous):

Find the value of k so that the line containing (4, -3k) and (2k, 5) has a slope of m=5/2.

OpenStudy (shamil98):

\[\frac{ y_2 - y_1 }{ x_2 - x_1 }\]

OpenStudy (anonymous):

Wouldn't the y's be 5 and -3k?, not five and four

OpenStudy (shamil98):

oh, oops, i mixed those up

OpenStudy (shamil98):

5 + 3k / 2k - 4 = 5/2

OpenStudy (shamil98):

5 + 3k = 5/2 (2k-4) 5 + 3k = 5(k-2) 5 + 3k = 5k -10 15 = 2k 15/2 = k